OpenStudy (anonymous):

. Is (x+y)2 equal to (x2+y2)? Explain why or why not using an example to support your answer.

OpenStudy (helder_edwin):

if x=y=2 then (x+y)^2=4^2=16 but x^2+y^2=2^2+2^2=4+4=8